Operation Summary for March 11, 2011 - March 14, 2011 |
|
The Integrated Luminosity for the period from 3/7/11 to 3/14/11 was 6.18 pb-1 (inverse picobarns). NuMI is off for target replacement. Friday March 11 The day shift began with the accelerator complex shutdown, due to ComEd power work, wrapping up. During the day shift, Operations reported a Main Injector (MI) scheduled power outage for MI-8 and the Kautz Road Substation. Operations reported that the MI recovery was prolonged due to MI-60 transductor issues. The Tevatron (TEV) came out of access at 3 PM. Operators searched and secured TEV sector E. Problems with an Antiproton Source (Pbar) power supply shunt (BS309) required a tunnel access to repair. During the evening shift, EE Support personnel replaced the MI triple transductor and its 300-amp transductor card. MI personnel found a closed beam valve (401), which prevented circulating beam. Operations reported that a MI RF station (MIRF16) bias power supply had tripped off. Operations found problems with a TEV higher order power supply (T:OD) and called in EE Support personnel to replace the power module. Pbar began stacking at 10 PM. Operators injected protons into the TEV near the end of the shift. Saturday March 12 The midnight shift began with the TEV in a dry squeeze, with Operations monitoring the Recycler stash (21.7E10) and the Pbar stack (8.2E10), and with the Linac, Booster, and MI operating. During the midnight shift, Operations reported that the TeV sector C4 BPMs had stopped responding at 1 AM. Operators tried multiple reboots and power cycles to no avail; Operations called in an Instrumentation technician. Operations reported that a Pbar Debuncher Septa (D:ISEPV) tripped off at 4 AM; operators investigated and reported a burned electronics odor around the power supply. Pbar personnel came in and replaced the power supply. Pbar resumed stacking at 5:15 AM. Stacking was interrupted twice by an AP0 dump cooling power supply trouble indication, which operators reset locally. The Instrumentation tech reported that he had the C4 BPM Front End (FE) working by 6:30 AM. D:ISEPV tripped off again at 6:30 AM; Pbar personnel came in, investigated, and started repairs. Operations reported that VSA MUX errors required operators to reset a VME (MI60S2) and to power cycle a MUX box. During the day shift, TEV personnel reported finding the TeV BPM signals at C4 reversed. Operations put the TEV into HEP shot setup at 2:10 PM and established store 8571 at 4 PM with an average initial luminosity (AIL) of 156E30. During the evening shift, LCW personnel accessed the MI to flush the Electron Cooling (ECool) magnets in order to remove a restriction in the LCW cooling lines. A Booster RF station (BRF18) tripped off; operators brought BRF19 online in its place. Sunday March 13 The midnight shift began with Operations monitoring store 8571 and the stash, and with Pbar stacking. During the midnight shift, Operations reported a sector B0 compressor room alarm at 12:30 AM. At 12:40 AM, store 8571 then quenched. Operations reported that a crack developed in the oil pressure regulator for the B0 Mycom cryo system compressor #2 and sprayed oil into the air and spilled onto the floor. The Fire Department contained the spill. Accelerator Division Safety and Cryogenics personnel came in and cleaned up the spill. Operations had the TEV recovered and with current on the bus by 7:35 AM. During the day shift, Operations put the TEV into shot setup at 10:05 AM and established store 8575 at 11:10 AM with an AIL of 237E30. During the evening shift, Operations terminated store 8575 at 6:05 PM due to Collider Detector at Fermilab (CDF) personnel needing to access their collision hall. D-Zero (D0) personnel also accessed their collision hall. Unfortunately, a Recycler kicker (KPS2A) pre-fired during a transfer of antiprotons and kicked 98E10 antiprotons out of the stash. Operations put the TEV into shot setup at 10:15 PM, but had to abandon the setup when operators found two bad CAMAC control cards for TeV corrector magnet power supplies; Operators replaced the control cards. Monday March 14 The midnight shift began with the TEV in shot setup, with Operations monitoring the stash, and with Pbar stacking. During the midnight shift, Operations put the TEV into shot setup at 1:45 AM and established store 8581 at 2:40 AM with an AIL of 310.2E30. Operations reported that BRF15 tripped off several times. The Plans for Monday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 03/14/2011 email Fermilab |
| Security, Privacy, Legal | |